# A Guide to Misinformation Detection Datasets

### v1

@@ 1 Introduction
We provide links to our unified dataset collection on HuggingFace, our code, and other outputs through our website. [^1] https://misinfo-datasets.complexdatalab.com/
Table 1: Characterizing 75 common misinformation datasets. Datasets are ordered by modality, then date, and topic.
Dataset Size Modality Topic Geographic region Language Time start Time end
AntiVax 15,465,687 Claims Health USA English 01/12/2021 31/07/2021
CoAID 301,177 Claims Covid-19 English 01/12/2019 01/09/2020
Counter-covid-19-misinformation 155,468 Claims Covid-19 International English 21/01/2020 20/05/2020
COVID-19-Rumor 7,179 Claims Covid-19 English 01/2020 03/2020
Covid-19-disinformation 16,000 Claims Covid-19 International 4 languages 01/2020 03/2021
Covid-vaccine-misinfo-MIC 5,952 Claims Covid-19 Brazil, Indonesia, Nigeria Eng., pt., id. 2020 2022
ESOC Covid-19 5,613 Claims Covid-19 International 35 languages 01/01/2020 1/12/2020
FakeCovid 7,623 Claims Covid-19 International 40 languages 04/01/2020 01/07/2020
FibVID 1,353 Claims Covid-19 International English 02/2020 01/2021
WICO 364,325 Claims Covid-19 English 17/01/2020 30/06/2021
Twitter16 818 Claims Various English 03/2015 12/2016
Rumors 1,022 Claims Various USA, UK, China, India English 01/05/2017 01/11/2017
ClaimsKG 74,066 Claims & Knowledge Graph Various International English 1996 2023
IFND 56,868 Claims & Images Various India English 2013 2021
Verite 1,001 Claims & Images Various International English 01/01/2001 01/01/2023
LIAR 12,836 Claims Politics USA English 2007 2016
LIAR-New 1,957 Claims Politics USA English & French 10/2021 11/2022
TruthSeeker2023 180,000 Claims Politics USA English 2009 2022
Check-COVID 1,504 Claims Covid-19 English
Climate-Fever 7,675 Claims Climate English
CMU-MisCOV19 4,573 Claims Covid-19 English
COVID-Fact 4,086 Claims Covid-19 English
FaVIQ 188,376 Claims Various English
FEVER 185,445 Claims Various USA English
FEVEROUS 87,026 Claims Various English
HoVer 26,171 Claims Various English
MediaEval 15,629 Claims & Images Various English, Spanish
MM-COVID 11,173 Claims Covid-19 International 6 languages
PHEME 62,445 Claims Newsworthy events International English
PubHealthTab 1,942 Claims Health North America English
Snopes Fact-news 4,550 Claims Various USA English
Twitter15 1,490 Claims Various English
X-Fact 31,189 Claims Various 25 languages
DeFaktS 105,855 Claims Various Germany German
MultiClaim 31,305 Claims Various International 39 languages
NLP4IF-2021 3,172 Claims Covid-19 International Ar., bul., eng.
Benjamin Political News 296 News articles Election USA English 02/2016 11/2016
BuzzFeedNews 2,282 News articles Election USA English 19/09/2016 27/09/2016
CT-FAN 2462 News articles Various Germany, USA, Canada English & German 2010 2022
Fake News Elections 38,333 News articles Politics USA English 04/2023 10/2023
FakeNews 486 News articles Politics USA English 01/2016 10/2017
FA-KES 804 News articles Syrian war Syria English 2011 2018
FANG-COVID 41,242 News articles Covid-19 Germany German 02/2020 03/2021
ISOT Fake News 44,898 News articles Politics International English 2016 2017
Italian disinformation 16,867 News articles & Tweets Election Italy English, Italian 01/01/2019 27/05/2019
Med-MMHL 40,601 News, tweets, images & LLM-generated Health USA English 01/01/2022 01/05/2023
NELA-GT-2020 1,779,127 News articles & Tweets Various USA English 01/01/2020 31/12/2020
ReCOVery 142,849 News articles & Tweets Covid-19 English 01/2020 05/2020
Spanish Fake News Corpus 572 News articles & Social media posts Various International Spanish 01/11/2020 31/03/2021
Weibo21 9,128 News articles Various China Chinese 12/2014 03/2021
BanFakeNews 50,000 News articles Various Bangladesh English
Celebrity 500 News articles Celebrities English
COVID-19 Fake News 10,700 News articles & Social media posts Covid-19 English
Fact-check-tweet 13,070 News articles & Tweets Various International 4 languages
FakeHealth 440,870 News articles & Social media posts Health USA English
FakeNewsAMT 480 News articles Various English
FakeNewsCorpus 9,408,908 News articles Various English
FakeNewsNet 23,196 News articles Politics & Celebrities English
FNC-1 49,972 News articles Various English
Misinfo Reaction Frames 25,100 News articles Global crises International English
MuMin 21,565,018 News articles, Tweets & Images Various International 41 languages
TI-CNN 20,015 News articles Politics USA English
BuzzFace 1,176,713 Social Media posts Election USA English 01/09/2016 30/09/2016
FacebookHoax 15,500 Social Media posts Hoaxes English 01/07/2016 31/12/2016
FACTOID 3,354,450 Social Media posts Politics USA English 01/2020 04/2021
Fakeddit 1,063,106 Social Media posts & Images Various English 19/03/2008 24/10/2019
VoterFraud2020 7,600,000 Social Media posts, Images & Videos Election USA English 23/10/2020 16/12/2020
MR2 14,700 Social Media posts & Images Rumor USA & China English & Chinese
Reddit 12,597 Social Media posts Various USA English
DBpedia 1,950,000 Wikipedia text Various International 14 languages
ICWSM 2,500 Images Election Brazil & India 10 languages 10/2018 06/2019
FaceForensics++ 1,800,000 Images Deepfakes
FCV-2018 380 Videos Various International 5 languages 2016 2017
Celeb-DF 5,639 Videos Celebrities
DEEPFAKETIMIT 640 Videos Various
2 Related Work
In recent years, the scientific community has shown a growing interest in fake news detection to mitigate the spread of false information. Within this evolving field, several surveys have emerged to offer comprehensive reviews and standardized evaluations. A pioneering effort by Shu et al. (2017) provided an early framework, defining fake news, detailing its characteristics, and summarizing detection techniques from a data mining perspective. Subsequent surveys, such as Oshikawa et al. (2018) and Zhou & Zafarani (2020), have explored alternative methodologies, focusing respectively on natural language processing (NLP) methods and interdisciplinary perspectives. However, while these surveys and others (Bondielli & Marcelloni, 2019; Gravanis et al., 2019) are valuable for providing a comprehensive overview of the state-of-art in fake news detection, they pay limited attention to existing datasets. Indeed, even if some emphasize the challenges of data collection or stress the importance of dataset quality, these surveys usually provide only superficial coverage of existing datasets, overlooking their specific content, details, and characteristics. Assessing the quality of these misinformation datasets is critical because they are often used to train and test models for misinformation detection and related tasks. A lack of quality data in this context implies that biases and erroneous conclusions could be introduced both in the development and in the validation process of these systems.
This gap has thus spurred the emergence of additional surveys dedicated to addressing these dataset-centric nuances, which can be categorized into two types. The first one focuses on categorizing existing datasets to guide the research community in their selection. For example, D’Ulizia et al. (2021) surveyed 27 datasets based on eleven characteristics (e.g., application purpose, type of disinformation, language, size, news content type, etc.) and compared these quantitatively. Another example, Sharma et al. (2019) summarized the characteristic features of 23 existing datasets, providing a clearer picture of those available to the public. However, these surveys have an important drawback; they often lack in-depth analysis. In fact, only descriptive characteristics are listed, thus neglecting key characteristics of their quality and effectiveness for future research. This is also the case for Ali et al. (2022) and Patra et al. (2022), which describe 26 and 7 datasets, respectively.
The second type of survey focuses on analyzing the quality, performance, and limitations of datasets. For instance, Abdali (2024) examines 10 datasets to identify some of these weaknesses and strengths. However, a broad approach is used to outline biases, which fails to detail the specifics of each dataset, leaving researchers uncertain about their individual quality. Another example, Hamed et al. (2023) highlight the limitations of 20 articles using publicly available datasets. While this approach provides a good overview of literature trends, a grey area remains regarding whether the errors in these 20 articles stem primarily from methodology or dataset issues. We also find the work of Pelrine et al. (2021), who evaluate the quality of six datasets, focusing on their potential spurious correlations with temporal information. Wu & Hooi (2022) expands the analysis of the spurious correlations issue to those induced by event-based collection, dataset merges, and labeling bias, using the Twitter15, Twitter16, and PHEME datasets. However, in both of these studies, the limited number of datasets analyzed fails to provide a comprehensive view of the diverse landscape of available datasets in this field. Similarly, Pelrine et al. (2023) highlights issues of ambiguous claims in the LIAR dataset, but does not expand their analysis beyond that one and their own LIAR-New dataset.
3 Survey
In this section, we introduce a collection of combined datasets on true and false information, including a total of 75 datasets, with 36 specifically focused on claims and statements. The full dataset contains a total of 120,901,495 observations, while the subset that we further analyze includes 1,612,933 observations. These data encompass a wide range of topics, including political issues, health concerns, and environmental questions, often related to the United States but also covering international news, headlines and online posts. The original labels within the datasets were assigned through a combination of expert evaluations and algorithmic methods. The following section provides a detailed summary of the data collection process and the characteristics of these datasets.
3.1 Collection Process
We then expanded our selection by rigorously examining the citations in scientific articles related to these initial datasets. Some articles listed available datasets in their literature reviews or surveys, enabling us to incorporate additional data. Through these combined approaches—which are further detailed in the Appendix—we identified 75 publicly accessible datasets, presented in Table 1.
For our analyses, we refined our selection to focus exclusively on datasets containing textual claims, defined here as short statements ranging from one to two sentences. Tweets are included in this definition, while lengthier online and social media posts are excluded. We chose to focus on this type of data because statements and claims are more concise than other forms of information, such as OP-ED and news articles, which often include opinions, commentary, and contextual details. This extraneous information can obscure the core claim or statement and introduce noise in the labeling process, as information can be partly true or false. Therefore, to test label quality, we initially focus on claims and statements to increase the reliability of our findings. This approach paves the way for future research to investigate the labeling of other types of content. Thus, with this filtering criteria, we reduced our selection to 36 datasets.
Table 2: Labelling approach and distribution for 36 claim datasets (subset of Table 1).
Dataset Labeling approach True claims (%) False claims (%) Other (%)
AntiVax Human expert 38.15 61.85 0
Check-COVID Human expert 37.92 16.48 24.92
ClaimsKG Human expert 19.83 70.58 9.59
Climate-Fever Human expert 52.64 16.48 30.88
CMU-MisCOV19 Human (N.S.) 17.75 74.74 7.51
CoAID Source-based categorization (T) & Human expert (F) 93.48 6.52 0
Counter-covid-19-misinformation Human (N.S.) 0 100 0
COVID-19-Rumor Human expert 26.16 51.27 22.57
Covid-19-disinformation Crowd-sourced 86.32 13.68 0
COVID-Fact Human-expert (T) & Algorithm-generated creation (F) 31.72 68.28 0
Covid-vaccine-misinfo-MIC Crowd-sourced 56.71 33.61 9.68
DeFaktS Human expert 58.86 41.14 0
ESOC Covid-19 Human expert 0 100 0
FakeCovid Human expert 0.86 94.99 4.15
FaVIQ Algorithm & Validation by human 49.77 50.23 0
FEVER Crowd-sourced 54.92 20.46 24.62
FEVEROUS Algorithm 57.77 38.77 3.46
FibVID Human expert 23.76 76.24 0
HoVer Crowd-sourced 58.74 41.26 0
IFND Human expert 66.65 33.35 0
LIAR Human expert 55.67 44.33 0
LIAR-New Human expert 15.02 84.98 0
MediaEval Source-based categorization 46.05 53.95 0
MM-COVID Human expert 71.74 28.26 0
MultiClaim Human expert 0 72.14 27.86
NLP4IF-2021 Crowd-sourced 95.15 4.85 0
PHEME Human expert and non-expert 22.14 77.86 0
PubHealthTab Human expert 52.47 23.79 23.74
Rumors Algorithm 13.78 70.47 15.76
Snopes Fact-news Human expert 16.93 63.57 19.50
TruthSeeker2023 Crowd-sourced 51.36 48.64 0
Twitter15 Human expert 50.07 24.83 25.10
Twitter16 Human expert 50.37 25.06 24.57
Verite Human expert 33.77 66.23 0
WICO Human expert 30.83 60.76 8.41
X-Fact Human expert 63.63 33.40 2.97
(T) indicates the method used to establish true claims
(F) indicates the method used to determine false claims
(N.S.) indicates that expertise is not specified
3.2 Claims Datasets
A summary description of each of our datasets can be found in Appendix A.2. Of these datasets, 12 consist of claims scraped from fact-checking or reliable websites, another 12 consist of tweets, and the remaining 11 comprise claims drawn from Twitter, the internet, social media, or news websites. There is variation in the topics of these datasets, but most focus on areas with significant societal impact where misinformation is prevalent and potentially harmful. For example, 16 of the datasets focus on health, vaccination, and COVID-19; 3 focus on political issues; 1 on environmental issues, and the rest covers various subjects, from culture, sport, the economy and so on. Unfortunately, a significant limitation of much of this data is the absence of information regarding the date the claim was made or fact-checked. This can potentially impact the accuracy of labeling, given that certain claims may have been true or false at the time they were made. In addition, this limitation affects the scope of our temporal leakage analysis. Consequently, scholars, and practitioners alike should be cautious when using these data.
4 Data Quality
Various factors affect the validity of the labels, which in turn impacts the overall quality of the data. In this section, we first discuss the strengths and weaknesses of each labeling technique used and discuss how we standardized the labels across studies. Next, we evaluate whether certain keywords are spuriously predictive of the veracity of the claims, and, finally, we examine if the datasets suffer from spurious temporal correlations.
4.1 Labeling Approach
The task of annotating statements is both crucial and challenging for anyone attempting to train a robust classifier for disinformation detection. Precise labeling is essential to ensure the classifier’s effectiveness, as it directly impacts its performance and reliability. Numerous approaches have been proposed in the literature to label true and false information. These approaches include expert and crowd-sourced annotation, source-based techniques, algorithmic methods, and a hybrid of these different approaches, all of which have been used in at least one of our 36 datasets (see Table 2). We describe these different approaches in turn to highlight their potential advantages and limitations.
Expert-based approach
Experts and fact-checkers are a small group of non-partisan professionals from various disciplines who manually verify the veracity of information. The result of these verifications are often published in fact-checking websites such as Politifact or Snopes. The strength of this approach lies in its rigorous review process, ensuring each piece of information is thoroughly evaluated, which leads to consistent reviews across fact-checkers. However, this method is not scalable and is costly (Zhou & Zafarani, 2020). As a result, experts must selectively choose the information they evaluate, which leads to many pieces of information going unchecked and potential biases in the selection of news and information that is evaluated (Lee et al., 2023; Markowitz et al., 2023; Walker & Gottfried, 2019).
Crowd-sourced approach
Crowdsourced fact-checking involves enlisting non-expert laypeople to assess the accuracy of online information. These evaluations are then aggregated to determine the veracity of the content. This approach is advantageous because it is more scalable, and laypeople can respond to misinformation much more quickly than professional fact-checkers (Zhao & Naaman, 2023). Additionally, this method has been shown to be effective in reducing the spread of misinformation and to produce veracity ratings similar to those of professional fact-checkers (Allen et al., 2021; Martel et al., 2024). However, crowdsourcing also has its limitations. It can be challenging to filter out evaluations from non-credible users and to ensure a balanced representation of users from different partisan backgrounds (Zhou & Zafarani, 2020; Martel et al., 2024).
Source-based approach
Source-based approaches to verifying information involve evaluating the domain or author of the content. Information is then rated as accurate if it comes from reliable sources and inaccurate otherwise. This method is more scalable than manual fact-checking, as it consists of evaluating the credibility of the source rather than each individual story. Additionally, this method is proven to be reliable, as experts generally rate news domains similarly (Lin et al., 2023). However, there are notable drawbacks. For instance, individual stories can vary in accuracy even within the same source, and not all content from low-quality outlets is necessarily false or misleading. Additionally, source familiarity significantly influences the perceived trustworthiness of content. Sources that are unfamiliar are often less trusted, which can lead to unfair negative evaluations of high-quality but lesser-known sources (Pennycook & Rand, 2019; Williams-Ceci et al., 2023)
Algorithmic methods
Finally, algorithmic methods can also be used to evaluate the veracity of content using NLP or other ML techniques (Zhou & Zafarani, 2020). For example, Covid-fact uses a BERT-based classifier, FaVIQ uses T5-3B, and Rumors uses an approach based on a social graph. These methods offer significant advantages in scalability, as they can process vast amounts of data quickly and efficiently, making them suitable for large-scale verification tasks. However, their accuracy can be questionable in many cases, ranging from struggles with nuanced or context-specific content (Boukouvalas & Shafer, 2024), issues with transfer and generalization (Huang et al., 2020; Pelrine et al., 2021; 2023), or just generically poor performance (e.g., even state-of-the-art methods often have below 70% accuracy compared to human labels (Zhang & Gao, 2023; Pelrine et al., 2023)). Thus, the quality of algorithmic labels is often dubious.
Our label unification approach
In this survey, we use the original labels from the studies to create a more consistent categorical variable across datasets. This allows us to test the accuracy of the veracity of the labels in each dataset. Specifically, we classify content as true, false, or unknown. Although this categorical variable is less precise than the detailed scales used to classify some of the claims in the data, we adopted this approach to ensure consistency across all datasets. Information that is mostly true is classified as “true”, while content that is mostly false or hyperpartisan is classified as “false”. This method aligns with existing research indicating that people respond similarly to content that is false or hyperpartisan (Ross et al., 2021; Pennycook et al., 2020). Finally, claims that could not be verified or were ambiguous, such as those partially true or false, were classified as “unknown”. The percent of true and false claims in each dataset using this coding scheme is shown in Table 2.
4.2 Spurious Keyword Correlations
We next evaluate whether there are certain keywords that overpredict misinformation in the claim datasets. We adapt the approach that Pelrine et al. (2021) used to check for spurious temporal correlations. Specifically, we trained a random forest classifier with the 40 most frequent words in each dataset, after removing stop-words and excluding datasets containing only tweet IDs (since in that case the claims are inaccessible), as well as those with single veracity labels that include only false statements. This analysis was conducted in two stages: first by incorporating only the labels True and False, and second, by also including the label Other. Utilizing scikit-learn, we set a maximum tree depth of 20 and retained the other default settings. The macro F1 results are shown in Table 3. We compare with a baseline of predicting randomly according to the class label distribution and no other information, providing a reference point for assessing the predictive power of the keywords. Thus, keywords in a dataset are significantly predictive if their performance significantly exceeds the baseline.
Table 3: Keywords correlations evaluation. A high predictivity score which far exceeds its corresponding baseline, means that the keywords provide an unrealistically strong prediction. All numbers are % macro F1.
Dataset Keywords Predictivity T-F Baseline T-F Keywords Predictivity T-F-O Baseline T-F-O
Check-COVID 44.9 47.3 22.9 31.7
ClaimsKG 43.8 50.0 27.6 32.6
Climate-Fever 44.9 49.5 23.0 33.7
CoAID 60.9 50.2 60.9 33.3
COVID-19-Rumor 48.9 50.7 37.5 32.2
COVID-FACT 40.6 52.3 - -
DeFaktS 37.3 50.3 - -
FakeCovid 50.0 50.0 33.3 49.9
FaVIQ 34.8 50.4 - -
FEVER 41.3 50.0 - -
FEVEROUS 37.4 49.5 24.4 32.6
FibVID 53.3 49.6 - -
HoVer 38.1 49.5 - -
IFND 82.2 50.0 - -
LIAR 44.7 49.8 - -
LIAR-New 46.0 52.3 - -
MM-COVID 77.1 51.2 - -
NLP4IF-2021 48.8 52.4 - -
PubHealthTab 42.6 49.5 31.0 33.5
Rumors 45.6 48.6 27.6 28.8
Snopes Fact-news 44.1 48.3 25.9 30.9
TruthSeeker2023 66.8 50.0 - -
Twitter15 62.2 48.9 35.7 33.7
Twitter16 66.4 43.3 55.1 38.6
Verite 39.9 49.0 - -
X-Fact 45.8 49.8 28.8 34.0
We particularly flag four datasets for spurious correlations between certain words and labels: IFND, MM-COVID, Twitter16, and TruthSeeker2023. For example, consider the latest, Truthseeker2023. Nearly all tweets here mentioning politicians are labeled as “false”, with only those containing “Trump” showing more variation (see also Appendix A.3). Obviously, in the real world, tweets mentioning politicians are obviously not exclusively false. Thus, models trained on data like Truthseeker2023 risk generalizing inaccurate results, and doing so on topics extremely sensitive to bias like discussion of politicians. Moreover, these findings extend beyond political names. Terms like “michigan”, “vaccines”, “immunity”, “pfizer”, and so on are consistently labeled as false, while words like “marijuana”, “wealth”, “terrorism”, “radical” and others are always associated with the true label (see also Figure  6 in the Appendix). Similar patterns are also observed in the other three datasets with highest keyword predictivity. Therefore, we urge caution about training and testing models on these datasets, especially text-focused ones.
4.3 Spurious Temporal Correlations
Pelrine et al. (2021) highlighted how collecting data of different classes at different times can make temporal information unrealistically predictive. For example, discussion of particular news events can become excessively correlated with veracity labels, leading to classifiers that rely on these events having artificially inflated performance, that will not generalize to the real world where veracity cannot be determined by past events alone.
Like in the preceding section, we follow their proposal for evaluating datasets for this limitation, training a random forest classifier. As feature, we use either the first three digits of the tweet ID (which contain time information) as in Pelrine et al. (2021) for Tweet datasets, or the date itself for datasets which include it. For the latter, we encode it as the integer number of days since the first date in the dataset. We exclude from this analysis datasets without either form of temporal information.
Table 4: Temporal correlations evaluation. A high score here means time—and information correlated with it—is unrealistically predictive.
Dataset Evaluation Type Temporal Predictivity (F1)
CoAID Date 48.3
FakeCovid Date 49.7
FibVID Date 62.2
LIAR-New Date 53.7
Rumors Date 69.1
X-Fact Date 63.2
AntiVax TweetID 46.8
CMU-MisCOV19 TweetID 45.6
Covid-19-disinformation TweetID 46.5
MediaEval TweetID 72.2
Twitter15 TweetID 83.2
Twitter16 TweetID 96.5
WICO TweetID 40.7
Results are shown in Table 4. We first note that our findings on Twitter15 and Twitter16 are similar to Pelrine et al. (2021), confirming these datasets have extreme issues with spurious correlations in temporal information. They should not be used without carefully and explicitly addressing this limitation. While not as severe, we also see that MediaEval and Rumors also suffer from some significant spurious temporal correlations, and caution is advised. The rest of the datasets have a substantially better temporal balance, with the temporal feature offering little better than random performance. However, we note that only a small fraction of the total datasets include dates, and recommend that future datasets add this important information.
4.4 Partisanship and Ideological Lean
We conducted analyses using large language models (LLMs) to assess the partisan and ideological lean of claims within various datasets. By providing distributions of these leans, we offer an approach for selecting datasets that better align with specific research goals. We also aim to determine if the veracity of a statement is spuriously correlated with its partisan lean, because such an association could in turn lead to models inaccurately assessing truthfulness based on that lean alone. Our findings indicate that a slightly higher proportion of true statements across the datasets are predicted to lean Democratic (12.12%) compared to those leaning Republican (9.5%). In contrast, a larger proportion of false statements are predicted to lean Republican (19.44%) than Democratic (9.36%). However, these proportions vary significantly across datasets, highlighting the importance of careful dataset selection. For a full table of results and further analysis, please see Appendix A.4.
5 Evaluation
5.1 Baseline Performance
Table 5: State-of-the-art GPT-4 baselines, with and without web search.

### v2

@@ 1. Introduction
We provide links to our unified dataset collection CDL-MD on HuggingFace, our tools CDL-DQA and other code, and other outputs through our website misinfo-datasets.complexdatalab.com.
Table 1. Dataset quality assessments. A denotes a dataset that passes the evaluation criterion. The horizontal line divides claim and paragraph datasets.
in Keyword or Temporal indicates supervised methods may learn significant spurious correlations of the respective type.
in Feasibility indicates over a quarter of the data may be impossible to assess for veracity at all.
Dataset Keywords Temporal Feasibility
AntiVax - -
Check-COVID -
ClaimsKG
Climate-Fever -
CMU-MisCOV19 - -
CoAID
COVID-19-Rumor -
Covid-19-disinformation - -
COVID-Fact -
DeFaktS
ESOC Covid-19 - -
FakeCovid
FaVIQ -
FEVER -
FEVEROUS -
FibVID
HoVer -
IFND -
LIAR -
LIAR-New
MediaEval - -
MM-COVID -
MultiClaim - -
NLP4IF-2021 -
PHEME - -
PubHealthTab -
Rumors
Snopes Fact-news -
TruthSeeker2023 -
Twitter15
Twitter16
Verite -
WICO - -
X-Fact
BanFakeNews
BenjaminPoliticalNews -
Celebrity -
CT-FAN -
FA-KES
FakeNewsAMT -
FakeNewsCorpus - -
ISOT Fake News
TI-CNN
Figure 1. An overview of the modalities, subject counts, and languages of datasets in CDL-MD. For more details on these and other attributes like size, geographic region, and date of the data, please refer to the comprehensive table in Appendix A.3.
2. CDL-MD
We curate a combined collection of 75 datasets, named CDL-MD (Complex Data Lab Misinfo Datasets). The collection includes 36 specifically focused on claims and statements (defined here as short texts ranging from one to two sentences), and 9 specifically focused on paragraphs (longer than two sentences). The full collection contains a total of 120,901,495 observations, while the subset that we further analyze includes 1,741,146 observations. These data encompass a wide range of topics, modalities, languages, and other characteristics (partially summarized in Figure 1, detailed in Appendix A).
Harmonizing labels across different datasets is a crucial step to ensure comparability of results and robustness of analyses. Since each study employs its own criteria for classifying veracity (see Appendix A.7.1), in our collection, we use the original labels from the studies to create a more consistent categorical variable across datasets. Specifically, we classify content as true, false, mixed or unknown. Information that is mostly true is classified as true, while content that is mostly false is classified as false. Finally, ambiguous claims, such as those partially true or false, were classified as mixed, while claims that were unproven, unrelated or contained no information about their veracity were coded as unknown.
An extensive treatment of the construction process of this collection, including how we identified the datasets, additional descriptive information for each one, expanded discussion of the labeling approach, and other discussion, is included in Appendix A.
3. Data Quality
Shortcut learning is a serious barrier to predictive systems working in the real world (Geirhos et al., 2020). In this section, we assess datasets’ potential for teaching algorithms spurious keyword and temporal correlations. We then assess whether the inputs are making sufficiently complete and unambiguous claims for it to be feasible to predict their veracity at all. We provide the automated tools that enable this analysis through our website as the Complex Data Lab Dataset Quality Assessment (CDL-DQA).
3.1. Spurious Keyword Correlations
We first evaluate whether there are certain keywords that overpredict veracity in the datasets. We adapt the approach that (Pelrine et al., 2021) used to check for spurious temporal correlations. Specifically, we trained a random forest classifier with the 40 most frequent words in each dataset, after removing stop-words and excluding datasets containing only tweet IDs (since in that case the claims are inaccessible), as well as those with single veracity labels that include only false statements. This analysis was conducted in two stages: first by incorporating only the labels True and False, and second, by also including the label Mixed. Utilizing scikit-learn, we set a maximum tree depth of 20 and retained the other default settings. We then compare with a baseline of predicting randomly according to the class label distribution and no other information, providing a reference point for assessing the predictive power of the keywords. Thus, keywords in a dataset are significantly predictive if their performance significantly exceeds the baseline. The macro F1 scores for the true and false labels are presented in Figure 2 and Table 6 provides the results that also incorporate mixed labels and baselines.
Figure 2. Keywords correlations evaluation. A high predictivity score that far exceeds the 50% baseline, indicated by the dashed line, means that the keywords provide an unrealistically strong prediction. Green bars indicate datasets that pass this check, while red bars represent ones with spurious correlations. All numbers are % macro F1.
We particularly flag six claims datasets for spurious correlations between certain words and labels, as well as six datasets containing paragraphs. Respectively, these are: CoAID, IFND, MM-COVID, TruthSeeker2023, Twitter15 and Twitter16; and BanFakeNews, BenjaminPoliticalNews, Celebrity, CT-FAN, ISOT Fake News and TI-CNN. For example, consider Truthseeker2023. Nearly all tweets here mentioning politicians are labeled as “false”, with only those containing “Trump” showing more variation (see also Appendix B.1). Obviously, in the real world, tweets mentioning politicians are not exclusively false. Thus, models trained on data like Truthseeker2023 risk generalizing inaccurate results, and doing so on topics extremely sensitive to bias like discussion of politicians. Moreover, these findings extend beyond political names. Terms like “michigan”, “vaccines”, “immunity”, “pfizer”, and so on are consistently labeled as false, while words like “marijuana”, “wealth”, “terrorism”, “radical” and others are always associated with the true label (see also Figure  12 in the Appendix). Similar patterns are also observed in the other datasets with highest keyword predictivity. Therefore, we urge caution about training and testing models on these datasets, especially text-focused models.
3.2. Spurious Temporal Correlations
Pelrine et al. (2021) highlighted how collecting data of different classes at different times can make temporal information unrealistically predictive. For example, discussion of particular news events can become excessively correlated with veracity labels, leading to artificially inflated performance for classifiers that rely on these events, that will not generalize to the real world where veracity cannot be determined by event or topic alone.
Like in the preceding section, we assess this limitation by training a random forest classifier. As feature, we use either the first three digits of the tweet ID (which contain time information) as in (Pelrine et al., 2021) for Tweet datasets, or the date itself for datasets which include it. For the latter, we encode it as the integer number of days since the first date in the dataset. We exclude from this analysis datasets without either form of temporal information.
Figure 3. Temporal correlation evaluation. A high score that far exceeds the 50% baseline means time—and information correlated with it—is unrealistically predictive. Green bars indicate datasets that pass this check, while red bars represent ones with spurious correlations. All numbers are % macro F1.
Results are shown in Figure 3 and Table 8. We first note that our findings on Twitter15 and Twitter16 are similar to (Pelrine et al., 2021), confirming these datasets have extreme issues with spurious correlations in temporal information. They should not be used without carefully and explicitly addressing this limitation. Similarly, the paragraph datasets BanFakeNews, IsotFakeNews, and TI-CNN show high levels of spurious temporal leakage. While not quite as severe, we also see that MediaEval and Rumors also suffer from some significant spurious temporal correlations, and caution is advised. The rest of the datasets have a substantially better temporal balance, with the temporal feature offering little better than random performance. However, we note that only a small fraction of the total datasets include dates, and recommend that future datasets add this important information.
3.3. Feasibility
If a claim is too ambiguous, it may be impossible and meaningless—or even misleading—to assess its veracity, irrespective of the power of one’s assessment system. For example, it is impossible to evaluate the veracity of the claim “The senator said the earth is flat” without knowing which senator is referred to. If a dataset contains too many such examples, it will be problematic to train and evaluate algorithms on it. (Pelrine et al., 2023) performed analysis on a limited, manual scale and found problems of this type in the LIAR dataset. To extend this analysis to the field at large, 8 human expert annotators each labeled examples from 29 datasets. We also developed an AI annotator that gives a comparable assessment, providing both additional evidence and scalability.
Our annotators categorized the claim text according to the following schema:
  • Feasible: The statement provides enough context for an AI to determine its truthfulness with certainty.
  • Feasible with web search: Some key information is missing, preventing an AI from determining the claim’s truthfulness without retrieving additional data online.
  • Not feasible: The statement is too vague or incomplete for a web search to provide sufficient evidence for verification.
This was done for 1230 human-annotated claims dataset examples (uniformly distributed for approximately 42 per dataset) with most annotated by at least two annotators, and 8700 AI-annotated claims examples plus 2700 paragraph examples (300 per dataset). Full human annotator instructions and sampling procedures, LLM prompting procedures, and other supplementary information are available in Appendix B.3.
Figure 4. Feasibility assessed by 8 human expert and 1 AI annotators, averaged over all datasets. Without an evidence retrieval system (e.g., web search), most data is not feasible to assess for veracity.
We first aggregate over datasets, presenting the individual results of our 9 human and AI annotators in Figure 4. Our results show universal agreement that without an evidence retrieval (search) system, at least half of claims in these 29 datasets cannot be validly assessed for veracity. This suggests that such methods are often being evaluated on impossible tasks, and there is a severe risk that evaluation comparing such methods is determining not the best generalizable predictor but the best shortcut learner. We recommend that any proposed methods in this category conduct stringent Evaluation Quality Assessment (see Section 5), and suggest that applications requiring even slightly out of distribution predictions should ground them in evidence.
Figure 5. Evaluation of claims dataset feasibility. Even with evidence retrieval, most datasets have a concerning proportion of infeasible data.
Figure 6. AI annotator evaluation of feasibility over paragraph datasets. These datasets often have higher feasibility proportions than claims datasets.
We next focus on systems that do have access to retrieval, particularly open web search, and assess which datasets will have strong feasibility and therefore represent strong options for training and evaluation. We propose 75% feasibility as a generous threshold, allowing for a moderate amount of noise in both our assessment and in applications relying on the datasets. We focus on the human annotator average (detailed in Appendix B.3), noting that the AI annotator has decent alignment with human assessment, and advantages in sample size and scalability, but is typically a bit more generous. Regardless, both forms of assessment yield similar conclusions. We note that for some applications, up to a quarter of examples in the evaluation data representing noise may be far too much—for example, comparisons between methods with margins of a couple percentage points. Nonetheless, we see in Figure 5 that most datasets do not even meet this standard.
This has 3 main implications. First, there is an urgent need for higher quality claims datasets. Second, supplementary information beyond claim text, such as claim dates, authors, and other additional context may help alleviate this problem. However, it is critical to assess how the specific information a predictive method processes impacts the feasibility of the data it is being trained and evaluated on, to confirm that performance margins between methods represent real progress and not progress in predicting noise. Thus, third, we propose that even in the more favorable retrieval-augmented setting, an Evaluation Quality Assessment (Section 5) is paramount.
We finally turn to paragraph datasets. We find (Figure 6) that they have a higher feasibility rate than claims datasets. This result makes sense considering that having a greater volume of text allows for more context, which reduces the amount of ambiguity in classifying misinformation. Thus, while it can be more complex to assess veracity with longer inputs, evaluation in this setting may be paradoxically easier. Since recent advances in LLMs (and multimodal systems that can process other types of inputs) provide many options for working with longer data, prioritizing this and other context-rich settings may be beneficial for the field.
4. Evaluation
4.1. Baseline Performance
Table 2. State-of-the-art GPT-4 baselines, with and without web search.